OpenStudy (anonymous):

One of two complementary angles is 20 degrees larger than the other. Find the measure of each Pleasee help :)))

OpenStudy (aripotta):

what do complementary angles add up to?

OpenStudy (unklerhaukus):

lets say the angles are \[\theta\qquad \& \qquad \phi\] complementary angles add to 90° so\[\theta+\phi=90°\] we also know that one of the angle is 20° larger than the other , so\[ \theta=\phi+20°\] ________ find the solution to the two equations simultaneously by substitution , sub theta from the second equation into the first \[\phi+20°+\phi=90°\] solve this for phi, back substitute to find theta,

OpenStudy (anonymous):

i have trouble doing simultaneous equations

OpenStudy (aripotta):

ok, all you have to do is (20 + x) + x = 90

OpenStudy (aripotta):

to find what x is

OpenStudy (aripotta):

so, combine like terms (x and x) first

OpenStudy (anonymous):

so.. it'll be 2x?

OpenStudy (aripotta):

yep. 20 + 2x = 90

OpenStudy (aripotta):

and now it's simple. just solve for x

OpenStudy (aripotta):

do you know how to solve for x?

OpenStudy (anonymous):

oh okay so it'll be 35?

OpenStudy (aripotta):

yea, so one angle is 35 degrees and the other is 35 + 20

OpenStudy (anonymous):

55?

OpenStudy (aripotta):

yep

OpenStudy (aripotta):

now double check by seeing if 35 + 55 = 90

OpenStudy (anonymous):

yeah it does thanks

OpenStudy (aripotta):

no problem :)